OpenStudy (anonymous):

(These questions are the ones i got wrong on my test I need someone to help me answer them please.) What are the solutions of the system? y=x^2-3x+2 y=4x-4 B. (1,20) and (6,0) C. (1,0) and (6,20) D. No solution

OpenStudy (anonymous):

Question?

OpenStudy (anonymous):

I just posted it

OpenStudy (cwrw238):

use substitution method as the terms in x both equal y, we have:- 4x - 4 = x^2 - 3x + 2 solve this for x

OpenStudy (anonymous):

can you help me work it out step by step

OpenStudy (cwrw238):

ok first subtract 4x from both sides -4 = x^2 -3x - 4x + 2 add 4 to both sides: 0 = x^2 -7x + 6 now we have a quadratic equation to solve can you do this?

OpenStudy (anonymous):

no I guess thats why i got most of them wrong

OpenStudy (cwrw238):

this will factor to a binomial of the form (x + a)(x + b) = 0 now x^2 - 7x + 6 = 0 to work out a and b we need 2 numbers whose product = last number 6 and whose sum = -7 look at possible factors of + 6: + 2 * +3 -2 * -3 +1 * +6 -1 * -6 which pair of factors will give us -7 when added?

OpenStudy (anonymous):

hold on one minute let me write all this down in my nots on how to do it

OpenStudy (anonymous):

*notes

OpenStudy (anonymous):

would -1*-7, +1*+7

OpenStudy (cwrw238):

it would be -1 and -6 ( because -1*-6 = +6 and -1 + -6 = -7) do you get that?

OpenStudy (anonymous):

yes i justed figured since it was on the one for 6 it wouldnt be here

OpenStudy (cwrw238):

perhaps the * confused you maybe i should have listed the factors as +2 and +3 etc see what i mean?

OpenStudy (anonymous):

yes so whats next

OpenStudy (cwrw238):

we 've found -1 and -6 so (x - 1)(x - 6) = 0 when two entities are multiplied to give 0 then either of them must = 0 , so either x-1 = 0 or x - 6 = 0 giving x = 1 or x = 6

OpenStudy (cwrw238):

the y values are obtained by plugging in x = 1 and x = 6 into y = 4x - 4 so one of them is y = 4(1) - 4 = 0 and solution is (1.0) the other one is 4(6) - 4 = ?

OpenStudy (anonymous):

would it be (6,20)?

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

Thanks so much

OpenStudy (cwrw238):

yw